Tickets for a concert were sold to adults for $3 and to students for $2. If the total receipt were $824 and twice as many adult tickets as student tickets were sold, then how many of each sold?
Answer by edjones(8007)

Let a be the number of adult tickets and s be the number of student tickets.
a=2s
3a+2s=824
3*2s+2s=824 substitute 2s for a
8s=824
s=103
a=206
.
Check:
3*206 + 2*103 = 824 true.
.
Ed